# Tree Ensembles for Contextual Bandits

### v2

@@ 3.4 Tree Ensemble Thompson Sampling
In Algorithm 2, we propose to estimate the uncertainty in the predicted reward given an arm’s context in the same way as in Algorithm 1. However, in line 18 of Algorithm 2, instead of constructing confidence bounds, we sample mean rewards from the resulting distributions (here, interpreted as posterior distributions). Hence, the main difference from TEUCB is found in the way in which the uncertainty is used to guide exploration. Due to its similarity with standard Thompson Sampling, we call this algorithm Tree Ensemble Thompson Sampling (TETS).
Algorithm 2 Tree Ensemble Thompson Sampling
1:  Input: Number of rounds $T$, number of initial random selection rounds $T_{I}$, number of trees in ensemble $N$, exploration factor $\nu$, tree ensemble regressor $F(\cdot\>;\{\cdot,\cdot\})$.
2:  for $t=1$ to $T_{I}$ do
3:     Randomly select and play an arm $a_{t}$
4:     Observe context $\mathbf{x}_{t,a_{t}}$ and reward $r_{t,a_{t}}$
5:  end for
6:  for $t=T_{I}+1$ to $T$ do
7:     Observe contexts $\{\mathbf{x}_{t,a}\}_{a=1}^{K}$
8:     for $a=1$ to $K$ do
9:        Initialize arm parameters:$\tilde{\mu}_{t,a}\leftarrow 0,\quad\tilde{\sigma}^{2}_{t,a}\leftarrow 0$
10:        for n = 1 to $N$ do
11:           Assign leaf values:
12:           $(o_{t,a,n},s_{t,a,n},c_{t,a,n})$$\leftarrow F_{n}\left(\mathbf{x}_{t,a};\left\{(\mathbf{x}_{i,a_{i}},r_{i,a_{i}})\right\}_{i=1}^{t-1}\right)$
13:           Estimate expected partial reward distribution: $\tilde{\mu}_{t,a,n}\leftarrow o_{t,a,n},\quad\tilde{\sigma}^{2}_{t,a,n}\leftarrow\frac{s^{2}_{t,a,n}}{c_{t,a,n}}$
14:           Increment arm parameters:
15:           $\tilde{\mu}_{t,a}\leftarrow\tilde{\mu}_{t,a}+\tilde{\mu}_{t,a,n}$
16:           $\tilde{\sigma}^{2}_{t,a}\leftarrow\tilde{\sigma}^{2}_{t,a}+\tilde{\sigma}^{2}_{t,a,n}$
17:        end for
18:        Sample reward $\tilde{r}_{t,a}\sim\mathcal{N}(\tilde{\mu}_{t,a},\nu^{2}\tilde{\sigma}^{2}_{t,a})$
19:     end for
20:     $a_{t}\leftarrow\text{argmax}_{a}\tilde{r}_{t,a}$
21:     Play $a_{t}$ and observe reward $r_{t,a_{t}}$
22:  end for
@@ 3.5 Extension to Combinatorial Bandits
Algorithm 1 and algorithm 2 are formulated to address the standard contextual MAB problem. However, TEUCB and TETS can easily be extended to the combinatorial semi-bandit setting. The main difference is in the way arms are selected, i.e., as super arms instead of individual arms. In Algorithm 1, this corresponds to modifying line 22 to
@@
The selected super arm $\mathcal{S}_{t}$ is subsequently played in line 23. Algorithm 2 can be similarly modified in lines 20 and 21, respectively. Furthermore, the set of observed context-reward pairs would include all rewards received for each base arm in the selected super arms individually, which are generally more than one per time step.
3.6 XGBoost Adaptation
@@
By using XGBoost regression models, we can extract $o$ directly from each individual leaf in the constructed ensemble. The sample variance $s^{2}$ is not accessible directly, but we can easily calculate it. As a GBDT method, XGBoost calculates its outputs as the average difference between the target value and the value predicted from the collection of preceding trees in the ensemble, multiplied with a learning rate. Hence, the trees do not have predetermined weights as in, e.g., random forests. Instead, the relative size of the contributions of an assigned leaf to the final prediction depends on the particular paths that are traversed through the other trees, which may be different for every sample we observe in a leaf. Therefore, we cannot estimate $s^{2}$ from the variance in target values directly.
@@
However, XGBoost is augmented with many useful features, one of them being staged predictions (XGBoost Developers, 2022a). This means that we can propagate the predicted outputs on sub-ensembles up to a certain tree on all data samples and cache these outputs. After recording these we include the next tree in the prediction as well, without having to start over. By utilizing this technique, $s^{2}$ is easily estimated from the previously observed samples assigned to a particular leaf. Furthermore, $c$ is simply the number of such samples. This gives us all the pieces of the puzzle needed to apply XGBoost to the TEUCB and TETS algorithms. In order to reduce bias when calculating $o$, $s^{2}$ and $c$, one may split the previously observed samples into two distinct data sets; one which is used for building the trees, and a second for the value estimations.
3.7 Random Forest Adaptation
@@
Similar to XGBoost, random forests can also be incorporated into TEUCB and TETS. This is even less complicated in the case of random forests as the trees in a random forest do not depend on the predictions made by any of the other trees. Therefore, $s^{2}$ can be directly estimated from the variance in the target values, taking the tree’s weight of $1/N$ into account. Hence, all the quantities of interest (i.e., $o$, $s^{2}$, and $c$) are available from the trees of a random forest by looking at which observed data points are assigned to the terminal nodes.
@@
As with XGBoost, one may consider dividing the previously observed samples into different data sets for fitting and value estimation. Alternatively, another way of handling bias when using random forests is to consider only the out-of-bag samples for computing $o$, $s^{2}$, and $c$. The random forest algorithm employs the concept of bagging—resampling with replacement—when building trees, which means roughly one third of the samples will be omitted in the construction of any particular tree. Therefore, random forests yield a natural way for providing an independent set of data points for model evaluation without limiting the size of the training data set.
@@ 4 Experiments
In this section, we evaluate TEUCB and TETS, and compare them against several best-known algorithms for solving the contextual bandit problem. For the implementations, we use the XGBoost library (Chen and Guestrin, 2016) to build gradient-boosted decision trees as our tree ensembles, as well as the version of random forest found in the scikit-learn library (Pedregosa et al., 2011). We first evaluate the algorithms on benchmark datasets. We then study them in the combinatorial contextual setting for solving the real-world problem of navigation over stochastic road networks in Luxembourg. The agent’s cumulative regret is calculated by comparing the path to an oracle agent, who knows the expected travel duration for each edge given the current time of day. All the algorithms are evaluated w.r.t. cumulative regret averaged over ten random seeds and we report the average results.
@@ 4.1 Setup
In these experiments, we set the time horizon to 10,000 time steps (except for the Mushroom dataset, where a horizon of 8,124 is sufficient to observe it entirely). For each of those time steps, the agents are presented with a feature vector $\mathbf{x}$ drawn randomly without replacement, which they encode for the different arms as described above. Subsequently, the agents predict the rewards for the individual arms according to the algorithms used. TEUCB (Algorithm 1) and TETS (Algorithm 2) are described in Section 3. The TreeBootstrap (Elmachtoub et al., 2017), NeuralUCB (Zhou et al., 2020), NeuralTS (Zhang et al., 2021), LinUCB (Li et al., 2010) and LinTS (Agrawal and Goyal, 2013) algorithms are implemented according to their respective references.
@@ 4.2 Contextual Bandits
For this experiment, the data is collected from the UCI machine learning repository (Kelly et al., n.d.), where an overview of the used selected datasets is given in Table 1. It should be noted that Magic (Bock, 2007), which is short for Magic Gamma Telescope, contains only numerical features. The same is true for Shuttle (Statlog (n.d.), Shuttle), which is also known as Statlog. All features of Mushroom (Mushroom, 1987) are categorical, and Adult (Becker and Kohavi, 1996) has a balanced distribution between numerical and categorical features.
@@ 4.2.1 Implementation
For all datasets in Table 1, the neural network agents use a network architecture of one hidden layer with 100 neurons. Regarding NeuralUCB, NeuralTS, LinUCB, and LinTS, we search the same sets of hyper-parameters as described by Zhang et al. (2021), who run these agents on the same datasets, and select the parameters with the best performance. One difference in our experiments is that we add a dropout probability of 0.2 when training the neural networks, which seems to have a positive impact on the performance. Furthermore, we let each of the networks train for 10 epochs.
For all tree ensemble bandits, we use XGBoost and random forest regressors with MSE loss and ensembles of 100 trees.
Initial test runs revealed that the agents are robust w.r.t. different choices of hyper-parameters for XGBoost (XGBoost Developers, 2022b), and consequently, we settle on the default values. The only exception is the maximum tree depth, which we set to $10$. The same maximum tree depth is used for the implementation of random forests. For the exploration factor of TEUCB and TETS, we select $\nu=1$. As for the decision tree (DT) algorithm used for TreeBootstrap, we use the scikit-learn library to employ CART (Breiman et al., 2017), which is free of tunable parameters. This is similar to the implementation presented in the original work on TreeBootstrap (Elmachtoub et al., 2017).
@@
In order to avoid unnecessary computations and speed up the runs, TEUCB and TETS do not build new tree ensembles from scratch at each time step. Instead, they only consider which leaves the latest observation assigned in each tree and update the corresponding $o$, $s^{2}$, and $c$ parameters. Initially, when new observed data points may have a relatively large effect on the optimal tree ensemble, re-building happens more frequently. However, as the effect that the data points are expected to have on the tree architectures degrades over time, less re-building takes place. More precisely, re-building happens when the function $\lceil 8\ln(t)\rceil$ increases by one compared to the previous time step.
@@
Table 2: Average regret accumulated by agents after the final step, with standard deviation, and number of hours required to run one experiment with CPU and GPU, respectively.
@@
TEUCB-XGBoost
Mean 1532.5 1685.7 69.2 171.4
SD $\pm$ 27.8 $\pm$ 27.7 $\pm$ 8.6 $\pm$ 40.0
@@
GPU - - - -
TETS-XGBoost
Mean 1548.0 1703.2 78.6 165.7
SD $\pm$ 30.9 $\pm$ 25.2 $\pm$ 11.8 $\pm$ 31.2
@@
GPU - - - -
TreeBootstrap-XGBoost
Mean 1693.2 1836.1 91.6 164.1
SD $\pm$ 18.8 $\pm$ 36.7 $\pm$ 5.6 $\pm$ 24.4
@@ 4.2.2 Experimental Results
The results of the experiments described above are presented in Fig. 1 and Table 2. We observe that the tree ensemble methods consistently outperform all other models by a large margin, both with XGBoost and random forest, and yield significantly lower cumulative regrets. Furthermore, TEUCB and TETS tend to perform better than TreeBootstrap on the adult, magic, and mushroom data sets, all of which contain only two arms.
On the shuttle data set, with seven different arms, the TreeBootstrap agents show comparable and even slightly better average performance in terms of regret minimization. It appears that TreeBootstrap’s assigning of a separate tree-based model for each arm is beneficial for this problem. However, this comes at the expense of having to fit multiple models in each time step, which is time-consuming and computationally demanding. Furthermore, its inability to generalize the reward predictions of distinct but related arms may in some cases be disadvantageous, especially for larger arm sets. Comparing TEUCB vs. TETS, and XGBoost vs. random forest, there is no clear winner as they all tend to perform comparable (and very effective) on different data sets. In addition to regret minimization, the CPU experiments indicate that TEUCB and TETS are significantly more efficient than their neural counterparts from a computational perspective. Notably, LinUCB, LinTS, and TreeBootstrap-DT are the most efficient methods in terms of computation. However, they fail to match the level of regret minimization as most other agents for our data sets.
@@ 4.3 Combinatorial Contextual Bandits
Figure 2: Experimental results on real-world road network navigation in Luxembourg. Fig. 2(b) shares the legend provided in Fig. 2(a).
@@ 4.3.1 Implementation
We model the road network of Luxembourg via graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, with $|\mathcal{V}|=2,247$ vertices and $|\mathcal{E}|=5,651$ edges. The vertices represent intersections in the road network, and the edges represent individual road segments connecting the intersections. In this scenario, edges correspond to base arms, and paths (i.e., ordered sequences of edges) correspond to super arms. Each vertex has a coordinate consisting of longitude, latitude, and altitude values. For all edges $e\in\mathcal{E}$, the contextual vector $\mathbf{x}_{e}$ describes each road segment in the network. The agent is presented with a vector containing contextual data according to Table 3.
Table 3: The variables included in the contextual vector describing each edge in the graph, presented to the agent when navigating in the road network.
@@
As the graph contains more than 5,000 edges, the agents need to learn how the contextual features impact the expected travel time. The road types in the graph are highways, arterial roads, and residential streets, with a total length of 955 km. In our experiments, we specifically study two problem instances (characterized by different start and end nodes), referred to as problem instance 1 and problem instance 2. The paths selected by the different agents during a single run are visualized for problem instance 1 in Fig. 3.
@@
For this experiment, the neural agents utilize a neural network with two hidden layers, both containing 100 fully connected neurons. A dropout probability of 0.2 is used and after a parameter search over the same sets of values as in Section 4.2.1, we set $\lambda=0.1$ and $\nu=0.001$. Furthermore, the network is trained over 10 epochs with the option of early stopping if the MSE loss does not tend to keep improving. We implement the tree ensemble bandits utilizing XGBoost and random forest regressors with maximum tree depths of $10$; all other hyper-parameters being set to their default values. The number of trees is set to $N=100$ for all agents with tree ensembles, and the initial random selection is set to $T_{I}=10$. An exploration factor of $\nu=1$ is being used for TEUCB and TETS and the frequency of re-building their tree ensembles is the same as described in Section 4.2.1.
Figure 3: The road network of Luxembourg shows the trajectories of different agents for the experiment on problem instance 1, where the corresponding cumulative regret is presented in Fig. 2(a).

### v3

@@ 3.4 Tree Ensemble Thompson Sampling
Using a TS-based approach, we propose to estimate the uncertainty in the predicted reward given an arm’s context in the same way as in the case with UCB. However, in line 23 of Algorithm 1, instead of constructing confidence bounds, we sample mean rewards from the resulting distributions (here, interpreted as posterior distributions). Hence, the main difference from TEUCB is how the uncertainty is used to guide exploration. Due to its similarity with standard Thompson Sampling, we call this algorithm Tree Ensemble Thompson Sampling (TETS).
@@ 3.5 Extension to Combinatorial Bandits
The framework outlined in Algorithm 1 is formulated to address the standard contextual MAB problem. However, TEUCB and TETS can easily be extended to the combinatorial semi-bandit setting. The main difference is in the way arms are selected, i.e., super arms instead of individual arms. In Algorithm 1, this corresponds to modifying lines 27 to
$\mathcal{S}_{t}\leftarrow\text{argmax}_{\mathcal{S}}\sum_{a\in\mathcal{S}}U_{t,a},$ (5)
and 29 to
@@
The selected super arm $\mathcal{S}_{t}$ at time $t$ is subsequently played in line 31. In this setting, the set of observed context-reward pairs would include all rewards received for each base arm in the selected super arms individually, which are generally more than one per time step.
3.6 Adaptation to XGBoost
Extreme gradient boosting (XGBoost) is a gradient boosting algorithm that utilizes ensembles of decision trees and techniques for computational efficiency (Chen and Guestrin, 2016). Gradient boosting creates new decision trees based on the gradient of the loss function. The ensemble output is based on the cumulative contribution of the decision trees within the ensemble.
@@ 3.6 Adaptation to XGBoost
When using XGBoost regression models, we can extract $o_{n}$ directly from the individual leaves in the constructed ensemble. The sample variance $s_{n}^{2}$ is not accessible directly, but we can easily calculate it. As a GBDT method, XGBoost calculates its outputs during the training procedure as the average difference between the target value and the value predicted from the collection of preceding trees in the ensemble, multiplied with a learning rate. Hence, the trees do not have predetermined weights as in, e.g., random forests. Instead, the relative size of the contributions of an assigned leaf to the final prediction depends on the particular paths that are traversed through the other trees, which may be different for every sample we observe in a leaf. Therefore, we cannot estimate $s_{n}^{2}$ from the variance in target values directly.
@@
However, XGBoost is augmented with many useful features, one of them being staged predictions (XGBoost Developers, 2022a). This means that we can propagate the predicted outputs on sub-ensembles up to a certain tree on all data samples and cache these outputs. After recording these we include the next tree in the prediction as well, without having to start over. By utilizing this technique, $s_{n}^{2}$ is easily estimated from the previously observed samples assigned to a particular leaf. Furthermore, $c_{n}$ is simply the number of such samples. This gives us all the pieces of the puzzle needed to apply XGBoost to the TEUCB and TETS algorithms. In order to reduce bias when calculating $o_{n}$, $s_{n}^{2}$ and $c_{n}$, one may split the previously observed samples into two distinct data sets; one which is used for building the trees, and a second for the value estimations.
3.7 Adaptation to Random Forest
Random forest (Ho, 1995) is a supervised machine learning algorithm that utilizes decision trees as base learners together with bootstrapped aggregating (bagging). Bagging models randomly sample from the training data with replacement—reducing the variance of the model by providing the base learners with diverse subsets from the complete dataset. In addition to bagging, a random forest randomly samples a subset of the features (Hastie et al., 2009). For classification tasks, majority voting is commonly used to make predictions, while in regression, the average or weighted average of predictions is a common method.
@@ 3.7 Adaptation to Random Forest
Similar to XGBoost, random forests can also be incorporated into TEUCB and TETS. This is even less complicated in the case of random forests as the trees in a random forest do not depend on the predictions made by any of the other trees. Therefore, $s_{n}^{2}$ can be directly estimated from the variance in the target values, taking the tree’s weight of $1/N$ into account. Hence, all the quantities of interest (i.e., $o_{n}$, $s_{n}^{2}$, and $c_{n}$) are available from the trees of a random forest by looking at which observed samples are assigned to the terminal nodes.
@@
As with XGBoost, one may consider dividing the previously observed samples into different subsets for fitting and value estimation. An alternative way for handling bias with random forests is to consider only the out-of-bag samples for computing $o_{n}$, $s_{n}^{2}$, and $c_{n}$. Random forests employ the concept of bagging—resampling with replacement—when building trees, which means roughly one-third of the samples will be omitted in the construction of any particular tree. Thus, they yield a natural way of providing an independent set of data samples for model evaluation without limiting the size of the training data set.
@@ 4 Experiments
In this section, we evaluate TEUCB and TETS and compare them against several well-known algorithms for solving the contextual bandit problem. For the implementations, we use the XGBoost library (Chen and Guestrin, 2016) to build gradient-boosted decision trees as our tree ensembles, as well as the version of random forests found in the scikit-learn library (Pedregosa et al., 2011). We first evaluate the algorithms on benchmark datasets. We then study them in the combinatorial contextual setting for solving the real-world problem of navigation over stochastic road networks in Luxembourg. The agent’s cumulative regret is calculated by comparing the path to an oracle agent, who knows the expected travel duration for each edge given the current time of day. All the algorithms are evaluated w.r.t. cumulative regret averaged over ten random seeds and we report the average results. The setup of the experiments is inspired by and adapted from Russo et al. (2018).
@@ 4.1 Setup
In these experiments, we set the time horizon to 10,000 (except for the Mushroom dataset, where a horizon of 8,124 is sufficient to observe it entirely). For each of those time steps, the agents are presented with a feature vector $\mathbf{x}$ drawn randomly without replacement, which they encode for the different arms as described above. Subsequently, the agents predict the rewards for the individual arms according to the algorithms used. TEUCB and TETS (Algorithm 1) are described in Section 3. The TreeBootstrap (Elmachtoub et al., 2017), NeuralUCB (Zhou et al., 2020), NeuralTS (Zhang et al., 2021), LinUCB (Li et al., 2010) and LinTS (Agrawal and Goyal, 2013) algorithms are implemented according to their respective references.
@@ 4.2 Contextual Bandits
For this experiment, the data is collected from the UCI machine learning repository (Kelly et al., n.d.), where an overview of the used selected datasets is given in Table 1. Magic (Bock, 2007), which is short for Magic Gamma Telescope, contains only numerical features. The same is true for Shuttle (Statlog (n.d.), Shuttle), also known as Statlog. All features of Mushroom (Mushroom, 1987) are categorical, and Adult (Becker and Kohavi, 1996) has a balanced distribution between numerical and categorical features.
@@ 4.2.1 Implementation
For all datasets in Table 1, the neural network agents use a network architecture of one hidden layer with 100 neurons. Regarding NeuralUCB, NeuralTS, LinUCB, and LinTS, we search the same sets of hyper-parameters as described by Zhang et al. (2021), who run these agents on the same datasets, and select the parameters with the best performance. One difference in our experiments is that we added a dropout probability of 0.2 when training the neural networks, which had a positive impact on the performance. Furthermore, each of the networks is trained for 10 epochs. For all tree ensemble bandits, we use XGBoost and random forest regressors with MSE loss and ensembles of 100 trees.
Initial test runs revealed that the agents are robust w.r.t. different choices of hyper-parameters for XGBoost (XGBoost Developers, 2022b), and consequently, we settle on the default values. The only exception is the maximum tree depth, which we set to $10$. The same maximum tree depth is used for the implementation of random forests. The number of initial random arm selections $T_{I}$ is set to 10 times the number of arms for all tree-based algorithms, and the exploration factor of TEUCB and TETS is set to $\nu=1$. As for the decision tree (DT) algorithm used for TreeBootstrap, we use the scikit-learn library to employ CART (Breiman et al., 2017), which is free of tunable parameters. This is similar to the implementation presented in the original work on TreeBootstrap (Elmachtoub et al., 2017).
@@
In order to avoid unnecessary computations and speed up the runs, TEUCB and TETS do not build new tree ensembles from scratch at each time step. Instead, they only consider which leaves the latest observation assigned in each tree and update the corresponding $o_{n}$, $s_{n}^{2}$, and $c_{n}$ parameters. Initially, when newly observed samples may have a relatively large effect on the optimal tree ensemble, re-building happens more frequently. However, as the effect that the samples are expected to have on the tree architectures degrades over time, less re-building takes place. More precisely, re-building happens when the function $\lceil 8\ln(t)\rceil$ increases by one compared to the previous time step.
@@
Table 2: Average regret accumulated by agents after the final step, with standard deviation, and number of hours required to run one experiment with CPU. Runtime in hours on GPU is also reported where it applies.
@@
TEUCB-XGBoost Mean $\pm$ SD 1532.5 $\pm$ 27.8 1685.7 $\pm$ 27.7 69.2 $\pm$ 8.6 171.4 $\pm$ 40.0
@@
TETS-XGBoost Mean $\pm$ SD 1548.0 $\pm$ 30.9 1703.2 $\pm$ 25.2 78.6 $\pm$ 11.8 165.7 $\pm$ 31.2
@@
TreeBootstrap-XGBoost Mean $\pm$ SD 1693.2 $\pm$ 18.8 1836.1 $\pm$ 36.7 91.6 $\pm$ 5.6 164.1 $\pm$ 24.4
@@ 4.2.2 Experimental Results
The results of the experiments described above are presented in Fig. 1 and Table 2. We observe that the tree ensemble methods consistently outperform all other models by a large margin, both with XGBoost and random forest, and yield significantly lower cumulative regrets. Furthermore, TEUCB and TETS tend to perform better than TreeBootstrap on the adult, magic, and mushroom data sets.
On the shuttle data set, with seven different arms, the TreeBootstrap agents exhibit comparable and even slightly better average performance in terms of regret minimization. It appears that TreeBootstrap’s assigning of a separate tree-based model for each arm is beneficial for this problem. However, this comes at the expense of having to fit multiple models in each time step, which is time-consuming and computationally demanding. Furthermore, its inability to generalize the reward predictions of distinct but related arms may in some cases be disadvantageous, especially for larger arm sets. Comparing TEUCB vs. TETS, and XGBoost vs. random forest (i.e., the different design choices within our framework), there is no clear winner as they all tend to perform comparably (and very effectively) on different data sets. In addition to regret minimization, the CPU experiments indicate that TEUCB and TETS are significantly more efficient than their neural counterparts from a computational perspective. Notably, LinUCB, LinTS, and TreeBootstrap-DT are the most efficient methods in terms of computation. However, they do not minimize regret as effectively as most other agents in our data sets.
@@ 4.3 Combinatorial Contextual Bandits
Figure 2: Experimental results on real-world road network navigation in Luxembourg.
@@ 4.3.1 Implementation
We model the road network of Luxembourg via a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, with $|\mathcal{V}|=2,247$ vertices and $|\mathcal{E}|=5,651$ edges. The vertices represent intersections in the road network, and the edges represent individual road segments connecting the intersections. In this scenario, edges correspond to base arms, and paths (i.e., ordered sequences of edges) correspond to super arms. Each vertex has a coordinate consisting of longitude, latitude, and altitude values. For all edges $e\in\mathcal{E}$, the contextual vector $\mathbf{x}_{e}$ describes each road segment in the network. The agent is presented with a vector containing contextual data according to Table 3.
Table 3: The variables included in the contextual vector describing each edge in the graph.
@@
As the graph contains more than 5,000 edges, the agents need to learn how the contextual features impact the expected travel time. The road types in the graph are highways, arterial roads, and residential streets, with a total length of 955 km. In our experiments, we specifically study two problem instances (characterized by different start and end nodes), referred to as problem instance 1 and problem instance 2. The paths selected by the different agents during a single run are visualized for problem instance 1 in Fig. 3 (in the Appendix).
@@
For this experiment, the neural agents utilize a neural network with two hidden layers, both containing 100 fully connected neurons. A dropout probability of 0.2 is used and after a parameter search over the same sets of values as in Section 4.2.1, we set $\lambda=0.1$ and $\nu=0.001$. Furthermore, the network is trained over 10 epochs with the option of early stopping if the MSE loss does not tend to keep improving. We implement the tree ensemble bandits utilizing XGBoost and random forest regressors with maximum tree depths of $10$; all other hyper-parameters being set to their default values. The number of trees is set to $N=100$ for all agents with tree ensembles, and the initial random selection is set to $T_{I}=10$ paths. An exploration factor of $\nu=1$ is used for TEUCB and TETS and the frequency of re-building their tree ensembles is the same as described in Section 4.2.1.